# Learning Chaotic Dynamics with Neuromorphic Network Dynamics

### v1

@@ Abstract
This study investigates how dynamical systems may be learned and modelled with a neuromorphic network which is itself a dynamical system. The neuromorphic network used in this study is based on a complex electrical circuit comprised of memristive elements that produce neuro–synaptic nonlinear responses to input electrical signals. To determine how computation may be performed using the physics of the underlying system, the neuromorphic network was simulated and evaluated on autonomous prediction of a multivariate chaotic time series, implemented with a reservoir computing framework. Through manipulating only input electrodes and voltages, optimal nonlinear dynamical responses were found when input voltages maximise the number of memristive components whose internal dynamics explore the entire dynamical range of the memristor model. Increasing the network coverage with the input electrodes was found to suppress other nonlinear responses that are less conducive to learning. These results provide valuable insights into how a practical neuromorphic network device can be optimised for learning complex dynamical systems using only external control parameters.
@@ I Introduction
Neuromorphic computing aims to achieve efficient computation by emulating the brain’s powerful information processing abilities [1, 2, 3, 4]. Two broad approaches that have been successfully demonstrated are: (i) electrical circuits that emulate point neuron spike models and/or spike–based learning models [5, 6]; and (ii) nano–electronic materials with inherent neuro–synaptic dynamics [1, 7, 8, 9]. The latter includes resistive switching memory (memristive) materials [10, 11, 12, 13]. Neuromorphic computing with memristors has been demonstrated with crossbar arrays, which can be used to perform physical matrix–vector operations to accelerate inference of artificial neural network models [14, 15]. Alternately, memristor crossbar arrays have also been used to implement reservoir computing (RC) [16]. Neuromorphic RC with memristive systems is a particularly promising approach given that electronic reservoirs with neuromorphic dynamics closely resemble the brain’s physical reservoir [17, 18]. However, a drawback of memristor crossbars for neuromorphic RC is their regular structure and limited scale.
@@
This study focuses on neuromorphic RC using memristive nanowire networks. These are complex, self-organised electrical circuits comprised of axon-like nanowires and synapse-like memristive junctions between overlapping nanowires, where conductance changes — the physical realisation of synaptic weights — are triggered by migration of atoms similar to the transmission of neurotransmitter molecules, constrained by physical equations of state and conservation laws [1]. Memristive nanowire networks exhibit emergent brain–like dynamics, such as dynamical phase transitions (continuous and discontinuous) and synchronisation. Previous studies have shown that different dynamical regimes may be exploited for learning [8, 19, 20, 21]. As a consequence of their self–organisation (i.e. their ability to acquire non–trivial structure–function properties), a single voltage pulse may have cascading effects that evolve all memristive elements in the neuromorphic network, in contrast to digital implementations of artificial neural networks, where large numbers of transistors must be individually addressed to update a similar number of network weights [2].
@@
Nanowire networks naturally self–organise into a Recurrent Neural Network (RNN) structure, with a functional connectivity that is highly suitable for RC. In standard algorithmic RC, the reservoir is typically a fixed random RNN of which each node evolves according to a mathematical activation function, and only a single linear output layer is trained with a computationally efficient linear regressor or classifier [22]. In this way, RC approaches can achieve excellent performance in various learning tasks involving temporal data [23, 24, 25], with training times that are orders of magnitude shorter than standard machine learning approaches. Physical RC, however, may operate differently because of the dynamics of the physical reservoir. In the case of neuromorphic RC with memristive nanowire networks, all the network dynamics are driven by the memristive edge junctions and are constrained by physics [26].
@@
To further investigate how memristive nanowire networks operate as physical reservoirs for neuromorphic RC, we present a simulation study focusing on a specific application: autonomous prediction of a multivariate chaotic time series, the well-known Lorenz63 system [27]. Prediction of the Lorenz63 system has been previously demonstrated with several different RC and related approaches [28, 29, 30, 31, 32, 33]. To the best of our knowledge, this study is the first to demonstrate autonomous prediction of the Lorenz63 attractor with neuromorphic RC. Understanding and predicting the behaviour of complex nonlinear dynamical systems is crucial for many scientific disciplines, including physics, climatology, neuroscience and biology, and for real–world applications such as financial markets and social networks [34, 35, 36]. Chaotic dynamical systems, in particular, are notoriously difficult to forecast due to their high sensitivity to initial conditions and highly nonlinear nature, resulting in aperiodic and categorically unpredictable dynamics. As such, the task of forecasting chaotic dynamics provides a challenging testing ground for exploring the limits of the capabilities of neuromorphic networks, and for testing the hypothesis that neuromorphic physical reservoirs can predict dynamical systems using the inherent physics of the neuromorphic system itself.
@@
This study primarily aims to gain deeper insights into the inner workings of memristive nanowire networks for learning via neuromorphic RC. No attempt is made to achieve state-of-the-art prediction results, such as those already achieved with random feature models[33], Long-Short Term Memory (LSTM) networks [37] or algorithmic RC models [38]. Rather our study provides the first proof of concept that analog devices consisting of a physical neuromorphic network is capable of performing learning tasks, through controlling the network via only the external input parameters. The forecast skill we obtain in our study for a network with $2,000$ nodes is far below those achieved by digital reservoir computers of the same size. However, as we will see performance increases with network size, and physical neuromorphic networks can easily be scaled up to millions of nodes [8, 39]; such network sizes are out of reach for digital reservoir computers. Also considering optimisations of the computational framework and the network itself (which is beyond the scope of this study), physical neuromorphic devices have great potential to outperform digital computers, with significantly faster and more energy efficient performance.
@@ II Methods
A neuromorphic reservoir was constructed using a physically–motivated model of a complex electrical circuit comprised of nonlinear resistive switching memory circuit elements known as memristors [1, 40, 41, 20, 19]. Given some input signal $\mathbf{u}(t)\in\mathbb{R}^{N_{\rm u}}$, the neuromorphic network is used to autonomously predict a dynamical time series $\mathbf{y}(t)\in\mathbb{R}^{N_{\rm y}}$, by computing the estimate $\mathbf{\hat{y}}(t)\in\mathbb{R}^{N_{\rm y}}$ of the teacher signal $\mathbf{y}(t)$. This is performed using a reservoir computing (RC) framework in which reservoir input values $\mathbf{r}_{\rm in}(t)$ are coupled to the input signal $\mathbf{u}(t)$ via a linear input layer with weight matrix $W_{\rm in}\in\mathbb{R}^{N_{\rm in}\times N_{\rm u}}$ and an input bias vector $\mathbf{b}_{\rm in}\in\mathbb{R}^{N_{\rm in}}$, such that
@@
with $W_{\rm in}$ and $\mathbf{b}_{\rm in}$ entries uniformly randomly sampled on intervals $[-w,w]$ and $[-b,-b/2]\cup[b/2,b]$, respectively (the choice of the bias sampling is explained in Sec. III.4). We choose $w=\alpha$, $b=5\alpha$ with $\alpha=0.2\,{\rm V}$. The $\alpha$ variable effectively serves as the input voltage scaling parameter which converts the input signals into appropriate voltages.
A graph representation of the neuromorphic network ${\cal N}$ consisting of $N$ nodes is abstracted as the reservoir (see Fig. 1), a high dimensional dynamical system which nonlinearly transforms a temporal signal $\mathbf{r}_{\rm in}(t)\in\mathbb{R}^{N_{\rm in}}$ into readout values $\mathbf{r}_{\rm out}(t)\in\mathbb{R}^{N_{\rm out}}$, with $N_{\rm out}\geq N_{\rm in}$. Only a subset $N_{\rm in}<N$ of all available nodes are used as input nodes, and $N_{\rm out}<N$ is the number of voltage readout nodes. In a machine learning context, these readouts $\mathbf{r}_{\rm out}(t)$ serve as high-dimensional dynamical feature embeddings that can be used to learn the nonlinear dynamics of the input data. The neuromorphic network model is described further in Sec. II.1.
@@
$W_{\rm out}=\operatorname*{{\rm arg\,min}}_{W_{\rm out}}\left(||W_{\rm out}\mathbf{r}_{\rm out}(t)-\Delta\mathbf{y}(t)||_{F}^{2}+\gamma||W_{\rm out}||^{2}_{F}\right),$ (3)
@@
$\frac{d\tilde{y}_{1}}{dt}=\rho_{1}(\tilde{y}_{2}-\tilde{y}_{1})\,,\quad\frac{d\tilde{y}_{2}}{dt}=\tilde{y}_{1}(\rho_{2}-\tilde{y}_{3})-\tilde{y}_{2}\,,\quad\frac{d\tilde{y}_{3}}{dt}=\tilde{y}_{1}\tilde{y}_{2}-\rho_{3}\tilde{y}_{3}\,.$ (5)
The constants were set to the standard values of $\rho_{1}=10$, $\rho_{2}=28$ and $\rho_{3}=8/3$ that generate chaotic dynamics. The normalised Lorenz values is recorded at equidistant sampling times $\Delta t=0.005$, and a training length of $135$ Lyapunov times was chosen to ensure a low training error (less than $0.01$ normalised root mean squared error). The sampling time $\Delta t$ is used as the discretisation time step of our network simulations. All three discretised variables of the Lorenz63 system in Eq. (5) were provided as input after being normalised to a mean of 00 and standard deviation of $1$, i.e. the $i$th element of $\mathbf{y}(t)$ is
@@
$t_{f}=\max_{t}{(\lambda_{\rm max}t|\mathcal{E}_{f}(t)\leq\theta)}\quad\text{ with }\quad\mathcal{E}_{f}(t)=\frac{||\mathbf{y}(t)-\mathbf{\hat{y}}(t)||^{2}}{\langle||\mathbf{y}-\langle\mathbf{y}\rangle||^{2}\rangle},$ (7)
where |||||$\cdot$||| | ⋅ | | denotes the Euclidean norm. A threshold of $\theta=0.4$ was chosen to compare results with related studies [28, 29]. The forecast time $t_{f}$ is measured in units of the Lyapunov time $\lambda_{\rm max}^{-1}$, where the largest Lyapunov exponent of the Lorenz63 system is $\lambda_{\rm max}=0.91$.
@@ II.1 Neuromorphic Network Model
The neuromorphic network $\cal N$ is comprised of a complex memristive circuit as a model of self–organised nanowire networks, which exhibit neuro–synaptic dynamics under electrical stimulation [1]. Due to its unique network topology, the neuromorphic networks were generated by modelling the bottom-up self-assembly process of the physical nanowire network; see Ref. [43] for detailed network generation and network topology analysis.
The neuromorphic network $\cal N$ is abstracted as a graph representation, with nodes representing nanowires and edges representing memristors. A network of $2,000$ nodes and $47,946$ edges, alongside another network of $500$ nodes and $9,905$ edges were used, with $5$% of all nodes serving as inputs and up to $90$% of all other nodes serving as readout nodes. Their graph representations are illustrated in Fig. 2.
Figure 2: Graph representations of simulated neuromorphic networks used in this study: (a) network with $2,000$ nodes and $47,946$ edges; (b) network with $500$-nodes and $9,905$ edges.
@@
The graph Laplacian of $\cal N$ is used to define the linear system of equations corresponding to Kirchhoff’s and Ohm’s laws for electrical circuits. For an arbitrary, non–trivial network with $N$ nanowires and $N_{\rm in}$ contact electrodes, for every specified time-step, the node voltages $\mathbf{r}$ and electrode currents $\mathbf{i}_{\rm in}$ are calculated by solving the following system of equations in response to the external input signals $\mathbf{r}_{\rm in}$
@@ II.2 Memristor Model
We use a threshold memristor model [19], given by the following expression for a single memristive junction with voltage $v(t)$ (indices omitted for clarity),
@@
The dimensionless state variable $x(t)$ describes the physico–chemical processes responsible for memristive switching, including electro–chemical metallisation and electron tunnelling. Here $q$ controls the decay of memory of the memristor, and $\eta$ controls the time scale of the memristor relative to the input signal.
@@
$g(x)=\left\{\begin{array}[]{ll}R_{\rm on}^{-1}&\quad|x|>x_{\rm crit},\\ R_{\rm off}^{-1}&\quad|x|\leq x_{\rm crit}.\\ \end{array}\right.$ (14)
@@
We remark that the nontrivial temporal evolution of the internal state variable in Eq. (10) is the only nonlinearity explicitly introduced in the neuromorphic network. If the state variables do not evolve (i.e. ${\dot{x}}_{ij}=0$ for all memristive junctions), the neuromorphic network reduces effectively to a network of linear resistors.
@@ II.3 Network Dynamics
Due to the memristive dynamics described above, the neuromorphic network exhibits coupled node–edge dynamics. An example of a network’s dynamical behaviour is presented in Fig. 4, which shows snapshots of conductance changes for different times as a voltage bias is applied to one node. As voltage is distributed to nodes according to their connectivity, edge conductances begin to evolve forming localised branches (Fig. 4(a)) that drive further node voltage redistribution, with the formation of the first high conductance path (Fig. 4(b), yellow) between the input and the ground node in the electrical circuit. As the input voltage persists, more parallel conductance paths form (Fig. 4(c)). These paths reset under reverse voltage polarity and/or via stochastic thermodynamic fluctuations (not modelled here to preserve interpretability of results).
@@
For arbitrary input voltage signals, the network is only dynamically active in certain parts at different times due to its connectivity. As shown in Fig. 4(d), network activity (defined by edge conductance changes when voltage difference are above a threshold set value $V_{\rm set}$) gradually increases, reaching a maximum of around $90$% before decreasing towards $50$%. After $0.5$ s, conductance changes slow down in tandem with reduced voltage differences between nodes. With multiple input nodes and more varied, bipolar input signals, the network’s behaviour follows similar dynamic patterns, but with richer complexity in its response. Recurrent edge–node interactions enable more nonlinear coupling between inputs and readouts, which may be advantageous for tasks involving multivariate time series with complex dynamics, such as the Lorenz system.

### v2

@@ Abstract
This study investigates how dynamical systems may be learned and modelled with a neuromorphic network which is itself a dynamical system. The neuromorphic network used in this study is based on a complex electrical circuit comprised of memristive elements that produce neuro–synaptic nonlinear responses to input electrical signals. To determine how computation may be performed using the physics of the underlying system, the neuromorphic network was simulated and evaluated on autonomous prediction of a multivariate chaotic time series, implemented with a reservoir computing framework. Through manipulating only input electrodes and voltages, optimal nonlinear dynamical responses were found when input voltages maximise the number of memristive components whose internal dynamics explore the entire dynamical range of the memristor model. Increasing the network coverage with the input electrodes was found to suppress other nonlinear responses that are less conducive to learning. These results provide valuable insights into how a physical neuromorphic network device can be feasibly optimised for learning complex dynamical systems using only external control parameters.
@@ I Introduction
Neuromorphic computing aims to achieve efficient computation by emulating the brain’s powerful information processing abilities [1, 2, 3, 4, 5]. Two broad approaches that have been successfully demonstrated are: (i) electrical circuits that emulate point neuron spike models and/or spike–based learning models [6, 2, 7]; and (ii) nano–electronic materials with inherent neuro–synaptic dynamics [8, 3, 9]. The latter includes resistive switching memory (memristive) materials [10, 11, 12, 13]. Neuromorphic computing with memristors has been demonstrated with crossbar arrays, which can be used to perform physical matrix–vector operations to accelerate inference of artificial neural network models [14, 15]. Alternately, memristor crossbar arrays have also been used to implement reservoir computing (RC) [16, 17]. Neuromorphic RC with memristive systems is a particularly promising approach given that electronic reservoirs with neuromorphic dynamics closely resemble the brain’s physical reservoir [18, 19]. However, a drawback of memristor crossbars for neuromorphic RC is their regular structure and limited scale [20, 21, 22, 23, 24, 25].
@@
This study focuses on neuromorphic RC using memristive nanowire networks. These are complex, self–organised electrical circuits comprised of axon-like nanowires and synapse-like memristive junctions between overlapping nanowires, where conductance changes — the physical realisation of synaptic weights — are triggered by migration of ions similar to the transmission of neurotransmitter molecules, constrained by physical equations of state and conservation laws [26]. Memristive nanowire networks exhibit emergent brain–like dynamics, such as dynamical phase transitions (continuous and discontinuous) and synchronisation. Previous studies have shown that different dynamical regimes may be exploited for learning [27, 28, 29, 30]. As a consequence of their self–organisation (i.e. their ability to acquire non–trivial structure–function properties), a single voltage pulse may have cascading effects that evolve all memristive elements in the neuromorphic network, in contrast to digital implementations of artificial neural networks, where large numbers of transistors must be individually addressed to update a similar number of network weights [31].
@@
Nanowire networks naturally self–organise into a Recurrent Neural Network (RNN) structure, with a functional connectivity that is highly suitable for RC. In standard algorithmic RC, the reservoir is typically a fixed random RNN of which each node evolves according to a mathematical activation function, and only a single linear output layer is trained with a computationally efficient linear regressor or classifier [32]. In this way, RC approaches can achieve excellent performance in various learning tasks involving temporal data [33, 34, 35], with training times that are orders of magnitude shorter than standard machine learning approaches. Physical RC, however, may operate differently because of the dynamics of the physical reservoir. In the case of neuromorphic RC with memristive nanowire networks, all the network dynamics are driven by the memristive edge junctions and are constrained by physics [36, 37, 38, 28, 27, 21, 39, 40, 41, 42].
@@
To further investigate how memristive nanowire networks operate as physical reservoirs for neuromorphic RC, we present a simulation study focusing on a specific application: autonomous prediction of a multivariate chaotic time series, the well–known Lorenz63 system [43]. Prediction of the Lorenz63 system has been previously demonstrated with several different RC and related approaches [44, 45, 46, 47, 48, 49, 50, 51, 52]. To the best of our knowledge, this study is the first to demonstrate autonomous prediction of the Lorenz63 attractor with physical neuromorphic RC. Understanding and predicting the behaviour of complex nonlinear dynamical systems is crucial for many scientific disciplines, including physics, climatology, neuroscience and biology, and for real–world applications such as financial markets and social networks [53, 54, 55]. Chaotic dynamical systems, in particular, are notoriously difficult to forecast due to their high sensitivity to initial conditions and highly nonlinear nature, resulting in aperiodic and categorically unpredictable dynamics. As such, the task of forecasting chaotic dynamics provides a challenging testing ground for exploring the limits of the capabilities of neuromorphic networks, and for testing the hypothesis that neuromorphic physical reservoirs can predict dynamical systems using the inherent physics of the neuromorphic system itself [56, 57, 58, 51, 31, 59].
@@
This study primarily aims to gain deeper insights into the inner workings of memristive nanowire networks for learning via neuromorphic RC. No attempt is made to achieve state-of-the-art prediction results for the Lorenz system, such as those already achieved with random feature models[49], Long-Short Term Memory (LSTM) networks [60] or algorithmic RC models [61]. Rather our study provides the first proof of concept that analog devices consisting of a physical neuromorphic network are capable of learning the Lorenz63 attractor, through controlling the network via only the external input parameters. The forecast performance achieved in our study for a network with $2,000$ nodes is far below those achieved by digital reservoir computers of the same size. However, as we show, performance increases with network size, and physical neuromorphic networks can readily scale to millions of nodes [62, 63]; such network sizes are out of reach for digital reservoir computers. Also considering optimisation of the computational framework and the network itself (which is beyond the scope of this study), physical neuromorphic devices have great potential to outperform digital computers, with significantly faster and more energy efficient performance [1, 3, 5].
@@ II Methods
A neuromorphic reservoir was constructed using a physically–motivated model of a complex electrical circuit comprised of nonlinear resistive switching memory circuit elements known as memristors [26, 37, 38, 28, 27]. Given some input signal $\mathbf{u}(t)\in\mathbb{R}^{N_{\rm u}}$, the neuromorphic network is used to autonomously predict a dynamical time series $\mathbf{y}(t)\in\mathbb{R}^{N_{\rm y}}$, by computing the estimate $\mathbf{\hat{y}}(t)\in\mathbb{R}^{N_{\rm y}}$ of $\mathbf{y}(t)$. This is performed using a reservoir computing (RC) framework in which reservoir input values $\mathbf{r}_{\rm in}(t)$ are coupled to the input signal $\mathbf{u}(t)$ via a linear input layer with weight matrix $W_{\rm in}\in\mathbb{R}^{N_{\rm in}\times N_{\rm u}}$ and an input bias vector $\mathbf{b}_{\rm in}\in\mathbb{R}^{N_{\rm in}}$, such that
@@
with $W_{\rm in}$ and $\mathbf{b}_{\rm in}$ entries uniformly randomly sampled on intervals [w,w][-w,w] and $b,-b/2]\cup[b/2,b]$, respectively (the choice of the bias sampling is explained in Sec. III.4). We choose $w=\alpha$, $b=5\alpha$ with $\alpha=0.2\,{\rm V}$. The $\alpha$ variable effectively serves as the input voltage scaling parameter which converts the input signals into appropriate voltages.
A graph representation of the neuromorphic network $\cal N$} consisting of $N$ nodes is abstracted as the reservoir (see Fig. 1), a high dimensional dynamical system which nonlinearly transforms a temporal signal $\mathbf{r}_{\rm in}(t)\in\mathbb{R}^{N_{\rm in}}$ into readout values $\mathbf{r}_{\rm out}(t)\in\mathbb{R}^{N_{\rm out}}$, with $N_{\rm out}\geq N_{\rm in}$. Only a subset $N_{\rm in}<N$ of all available nodes are used as input nodes, and $N_{\rm out}<N$ is the number of voltage readout nodes. In a machine learning context, these readouts $\mathbf{r}_{\rm out}(t)$ serve as high-dimensional dynamical feature embeddings that can be used to learn the nonlinear dynamics of the input data [64]. The neuromorphic network model is described further in Sec. II.1.
@@
[…]
@@
$\begin{split}\frac{d\tilde{y}_{1}}{dt}&=\rho_{1}(\tilde{y}_{2}-\tilde{y}_{1}),\\ \frac{d\tilde{y}_{2}}{dt}&=\tilde{y}_{1}(\rho_{2}-\tilde{y}_{3})-\tilde{y}_{2},\\ \frac{d\tilde{y}_{3}}{dt}&=\tilde{y}_{1}\tilde{y}_{2}-\rho_{3}\tilde{y}_{3}.\end{split}$ (5)
The constants were set to the standard values of $\rho_{1}=10$, $\rho_{2}=28$ and $\rho_{3}=8/3$ that generate chaotic dynamics. The normalised Lorenz values is recorded at equidistant sampling times $\Delta t=0.005$, and a training length of $135$ Lyapunov times was chosen to ensure a low training error (less than $0.01$ normalised root mean square error). The sampling time $\Delta t$ is used as the discretisation time step of our network simulations. All three discretised variables of the Lorenz63 system in Eq. (5) were provided as input after being normalised to a mean of 0 and standard deviation of $1$, i.e. the $i$th element of $\mathbf{y}(t)$ is
@@
where $\langle\cdot\rangle$ denotes the time average, and $\sigma_{\tilde{y}_{i}}$ is the standard deviation (across all training time) of the $i$th component of the Lorenz system.
@@
where $||\cdot||$ denotes the Euclidean norm. A threshold of $\theta=0.4$ was chosen to be comparable with related studies [44, 45]. The forecast time $t_{f}$ is measured in units of the Lyapunov time $\lambda_{\rm max}^{-1}$, where the largest Lyapunov exponent of the Lorenz63 system is $\lambda_{\rm max}=0.91$.
@@ II.1 Neuromorphic Network Model
The neuromorphic network $\cal N$ is comprised of a complex memristive circuit as a model of self–organised nanowire networks on a 2D plane, which exhibit neuro–synaptic dynamics under electrical stimulation [26]. Due to its unique network topology, the neuromorphic networks were generated by modelling the bottom–up self–assembly process of the physical nanowire network; see Ref. [66] for detailed network generation and network topology analysis.
The neuromorphic network $\cal N$ is abstracted as a graph representation, with nodes representing nanowires and edges representing memristors. The resulting network is undirected and connected (no isolated subgraphs). A network of $2,000$ nodes and $47,946$ edges, alongside another network of $500$ nodes and $9,905$ edges were used, with 55% of all nodes serving as inputs and up to 9090% of all other nodes serving as readout nodes. Their graph representations are illustrated in Fig. 2. Both networks have an average degree of $\simeq 40$, and a negatively skewed log-normal degree distribution (see Fig. 20 in Appendix C).
Figure 2: Graph representations of simulated neuromorphic networks used in this study: (a) network with $2,000$ nodes and $47,946$ edges; (b) network with $500$ nodes and $9,905$ edges. The corresponding degree distributions are shown in Appendix Fig. 20.
@@
The graph Laplacian of $\cal N$ is used to define the linear system of equations corresponding to Kirchhoff’s and Ohm’s laws for electrical circuits. For an arbitrary, non–trivial network with $N$ nanowires and $N_{\rm in}$ contact electrodes, for every specified time-step, the node voltages $\mathbf{r}$ and electrode currents $\mathbf{i}_{\rm in}$ are calculated by solving the following system of equations in response to the external input signals $\mathbf{r}_{\rm in}$ [67, 68]
@@ II.2 Memristor Model
We use a threshold memristor model [27, 37, 28], given by the following expression for a single memristive junction with voltage $v(t)$ (indices omitted for clarity),
@@
The physico-chemical processes responsible for memristive switching include electro-chemical metallisation (ECM) and electron tunnelling. Here, only electron tunnelling is explicitly modelled to update the dimensionless state variable $x(t)$. Other models have been proposed that capture ECM, but not electron tunnelling [69, 70]. Constant $q$ controls the decay of memory of the memristor, and $\eta$ controls the time scale of the memristor relative to the input signal.
@@
[…]
@@
We remark that the nontrivial temporal evolution of the internal state variable in Eq. (10) parametrises internal nonlinear conductance changes via Eq. (11). If the state variables do not evolve (i.e. ${\dot{x}}_{ij}=0$ for all memristive junctions), the neuromorphic network reduces effectively to a network of linear resistors.
@@ II.3 Network Dynamics
The neuromorphic network node dynamics are driven by the memristive dynamics described above. An example of a network’s dynamical behaviour is presented in Fig. 4, which shows snapshots of conductance changes for different times as a voltage bias is applied to one node. As voltage is distributed to nodes according to their connectivity, edge conductances begin to evolve forming localised branches (Fig. 4(a)) that drive further node voltage redistribution, with the formation of the first high conductance path (Fig. 4(b), yellow) between the input and the ground node in the electrical circuit. As the input voltage persists, more parallel conductance paths form (Fig. 4(c)). These paths reset under reverse voltage polarity and/or via stochastic fluctuations associated with atomic nanofilament formation and decay [72], which can also affect the path setting process. We do not model stochastic fluctuations here to preserve interpretability of results.
@@
For arbitrary input voltage signals, the network is only dynamically active in certain parts at different times due to its connectivity. As shown in Fig. 4(d), network activity (defined by edge conductance changes when voltage difference are above a threshold set value $V_{\rm set}$) gradually increases, reaching a maximum of around 9090% before decreasing towards 5050%. After $0.5$ s, conductance changes slow down in tandem with reduced voltage differences between nodes. With multiple input nodes and more varied, bipolar input signals, the network’s behaviour follows similar dynamic patterns, but with richer complexity in its response. Recurrent edge–node interactions enable more nonlinear coupling between inputs and readouts, which may be advantageous for tasks involving multivariate time series with complex dynamics, such as the Lorenz system.